### v2

# Tailoring Dynamical Codes for Biased Noise: The X3Z3 Floquet Code

Abstract
We propose the X3Z3 Floquet code, a type of dynamical code with improved performance under biased noise compared to other Floquet codes. The enhanced performance is attributed to a simplified decoding problem resulting from a persistent symmetry under infinitely biased noise, which surprisingly exists in a code without constant stabilisers. Even if such a symmetry is allowed, we prove that a general dynamical code with two-qubit parity measurements cannot admit one-dimensional decoding graphs, a key feature resulting in the high performance of bias-tailored stabiliser codes. Despite this limitation, we demonstrate through our comprehensive numerical simulations that the symmetry of the X3Z3 Floquet code renders its performance under biased noise far better than several leading Floquet code candidates. Furthermore, to maintain high-performance implementation in hardwares without native two-qubit parity measurements, we introduce ancilla-assisted bias-preserving parity measurement circuits. Our work establishes the X3Z3 code as a prime quantum error-correcting code candidate, particularly for devices with reduced connectivity, such as the honeycomb and heavy-hexagonal architectures.
I Introduction
Quantum error correction (QEC) [1, 2, 3, 4] should be understood as occurring both in space and time [5]. Taking advantage of the temporal dimension, Floquet codes [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20], or more generally dynamical codes [21, 22, 23], form a large class of error-correcting codes, which can achieve competitive fault-tolerant performance [6, 8, 9, 10, 13, 12, 11, 14, 16, 15, 17] while reducing the weights of check measurements performed for the error correction. Several of these codes [7, 9, 18, 13] also benefit from being defined on a lattice with sparser connectivity than that for the surface code [24, 25]: each qubit is only connected to three other qubits. In architectures where two-qubit parity check measurements are native, Floquet codes could achieve higher thresholds [10, 8, 17] and lower space-time overheads [8, 10] than the surface code. Without requiring additional connectivities, this code can be deformed around defective components caused by highly noisy qubits or gates [15, 26]. Moreover, dynamical codes allow for implementations of arbitrary Clifford, and even some non-Clifford gates, through low-weight parity check measurements [21]. In addition, dynamical measurement schedules can result in certain errors being self-corrected [20].
Although Floquet codes (in particular, the honeycomb code) have been studied under various noise models [9, 27], there have not been any Floquet codes that are specifically tailored for an improved performance under biased noise. A biased noise model is one in which a specific type of error, for example, phase errors, occurs more frequently than other errors, such as bit flip errors. This biased noise is typical to most quantum platforms, for example, superconducting fluxonium qubits [28], bosonic “cat” qubits [29, 30], spin-optical [31, 17], neutral atoms [32], quantum-dot spin qubits [33, 34] and Majorana qubits [35, 36].
For enhanced performance, quantum-error-correcting codes need to be designed such that they possess symmetries that can be utilised to simplify the error-syndrome decoding problem given the noise structure [37, 38, 39, 40, 41, 42, 43, 44, 45]. While there have been a number of proposals for bias-tailored static codes [38, 39, 41, 45, 46, 40, 47, 41, 44, 42, 48, 49, 43, 50, 51], designing Floquet codes for high performance under biased noise is still an open problem. Owing to the experimental relevance of biased noise and given the ease of implementation of Floquet codes which require only two-qubit parity measurements, it is therefore imperative to tailor Floquet codes for biased noise and study how the performance of such dynamical codes can be improved.
In this paper, we present the X3Z3 bias-tailored Floquet code, a Clifford-deformed [45] version of the Calderbank–Shor–Steane (CSS) Floquet code [18, 13]. Despite not having a fixed stabiliser group (as static codes have), the X3Z3 Floquet code surprisingly still possesses a symmetry under infinitely biased noise that allows for a simplified decoding in biased noise regimes. We perform an in-depth study of this code, along with the CSS Floquet code [18, 13], and two types of honeycomb codes: one introduced by Hastings and Haah [6] and the other by Gidney et al. [8]. We simulate all codes under biased-noise models, and find that the X3Z3 code has the best performance. Using a matching decoder, we find that, as the noise changes from fully depolarising to pure dephasing, the X3Z3 Floquet code threshold increases from $1.13\%$ to $3.09\%$ under a code-capacity noise model and increases from $0.76\%$ to $1.08\%$ under a circuit-level noise model mimicking hardwares with noisy direct entangling measurements. Furthermore, we show that its sub-threshold performance is also substantially better under biased noise than other Floquet codes.
Compared to its static counterparts, the X3Z3 Floquet code has an advantage that it can be realised using only two-qubit parity check measurements. This makes it particularly suitable for devices with constrained connectivities, such as the honeycomb and heavy-hexagonal lattice (currently IBM’s preferred superconducting-qubit architecture) [52, 53]. Moreover, we demonstrate that the two-qubit parity measurements of the Floquet code can be performed in a bias-preserving way even in hardware without direct entangling measurements, thus enabling high performance implementation in such devices.
To explain the performance of our bias-tailored Floquet code, we argue that other dynamical codes defined on the same architecture, and built from two-body measurements, would likely not have drastically improved performance compared to the X3Z3 Floquet code. To support this argument, we prove that decoding graphs of such dynamical codes under infinitely biased noise have connectivities that are too high for the decoding problem to be reduced to a simple decoding of repetition codes, as is the case for static codes [41, 45]. This can be understood as resulting from the fact that error syndromes of dynamical codes possess less symmetry than their static code counterparts.
The paper is laid out as follows. In Sec. II, we review the basics of Floquet codes together with two widely studied examples: honeycomb and CSS Floquet codes. Readers who are already familiar with Floquet codes can skip directly to Sec. III where we discuss our X3Z3 Floquet code which is obtained by bias-tailoring the CSS Floquet code. Crucially, we show that there exists a persistent symmetry in the code’s error syndrome under pure dephasing noise that allows for a simplified decoding. Subsequently, we introduce ancilla-assisted bias-preserving parity measurement circuits that allow for high-performance code implementation in devices without native entangling measurements. In Sec. IV, we present our simulation methods and noise models. In Sec. V, we then provide our simulation results for all of the four Floquet codes studied. To explain the code performance, in Sec. VI we prove that a class of dynamical codes on the honeycomb lattice has high-connectivity decoding graphs. We also prove that dynamical codes with only two-qubit parity check measurements cannot have repetition-code-like decoding graphs, as is the case with static codes. Finally, we conclude and present future research directions in Sec. VII. In the Appendices, we give a more detailed review of the basics of honeycomb and CSS Floquet codes, and present the details of our noise models, numerical simulations, and theorems.
II Background
II.1 Dynamical and Floquet codes
Figure 1: (a) CSS and (b,c) X3Z3 Floquet codes. (a,b) Left: Hexagonal lattice with qubits sitting on vertices and opposite boundaries identified. Plaquettes and edges are assigned one of three colours [red, green, blue; examples are shown in (a)] and one of two types (X- or Z-type for CSS and A- or B-type for X3Z3). (a) Right: CSS Floquet code measurement schedule. (b) Left: The X3Z3 Floquet code is obtained from the CSS code via Hadamard gates applied to shaded domains’ qubits. Right: Plaquette and edge (check) operators are A- or B-type, depending on their support basis on the shaded/unshaded domains. (c) The X3Z3 Floquet code measurement schedule. Arrows indicate the type and colour of the edge operator measured at each step, where the edges just measured (members of the ISG) are highlighted in the lattice. Uncoloured plaquettes host only a single type of stabiliser, either A or B, indicated by the letters in the plaquettes, while coloured plaquettes host both A- and B-type stabilisers in the ISG. One set of anti-commuting logical operators is shown by yellow and brown strings, where their $X$ ($Z$) bases on the qubit supports are highlighted in red (blue). The other set (not shown) is similar to the set shown but offset by three measurement subrounds and with an $X\leftrightarrow Z$ interchange of the qubit supports’ bases.
We begin by defining dynamical and Floquet codes. Here we consider the Floquet codes to be defined on the lattice of a two-dimensional colour code, which is trivalent and three-colourable. A trivalent lattice has each vertex incident to three edges, and a three-colourable lattice has every face assigned one of three colours in such a way that there are no two adjacent faces of the same colour. Throughout this paper, we will use the honeycomb lattice as an example of such a lattice (see Fig. 1).
We define qubits as residing on the vertices of the lattice and error-correction will be performed by measuring two-qubit operators defined on edges of the lattice (i.e., acting on qubits incident to a given edge). Each edge is coloured the same as the plaquettes it connects [see e.g., the coloured edges in the middle right of the lattice in Fig. 1(a)]. We perform error-correction by measuring edge operators in a particular sequence. After any given subround of edge measurements, the system will be in the $+1$-eigenstate of the operators in an “instantaneous stabiliser group” (ISG), which will change at each time step. The ISG at time step $t+1$ is defined as
$\begin{split}\text{ISG}_{t+1}=\langle S,\pm M\;|\;M\in\mathcal{M}_{t+1},\;S\in\text{ISG}_{t},\\ \text{such that }[S,M]=0,\;\forall M\in\mathcal{M}_{t+1}\rangle.\end{split}$ (1)
In the above, $\mathcal{M}_{t+1}$ is the set of “check” measurements performed at time step $t+1$. The ISGt+1 also includes “plaquette stabiliser operators” $S\in\text{ISG}_{t}$ which commute with all $M\in\mathcal{M}_{t+1}$. The check measurement operators are chosen in such a way that those check operators at time $t$ that have overlapping qubit supports with check operators at $t+1$ anti-commute. For Floquet codes, the measurement sequence is periodic, such that $\mathcal{M}_{t+T}=\mathcal{M}_{t}$ for some integer $T$. For such a code, we will be performing quantum memory experiments with $mT$ time steps, for some integer $m$. We will refer to $m$ as the number of “QEC rounds” in the experiment, while we will refer to $mT$ as the number of “measurement subrounds” in the experiment.
The logical operators at time $t$ are given by $\mathcal{C}(\mathrm{ISG}_{t})/\text{ISG}_{t}$, where $\mathcal{C}(\text{ISG}_{t})$ is the centraliser of ISGt, i.e., the group of Pauli operators commuting with all $S\in\text{ISG}_{t}$. A (potentially trivial) logical operator “representative” is some member of $\mathcal{C}(\text{ISG}_{t})$. Each nontrivial logical operator representative for the codes considered will be a string-like Pauli operator at each time step [see Fig. 1(c)]. To avoid anti-commuting with the next-subround edge measurements, certain check measurement results along a logical operator’s path have to multiplied into that logical operator. Hence, the logical operators will evolve from one time step to the next.
We can detect errors if we can find sets of measurements, called detectors, that always multiply to the value $+1$ in the absence of noise, thus registering no error. Over some number of QEC rounds we will have extracted several detector outcomes. A detector (or decoding) hypergraph is formed by first defining a node for each (independent) detector in the code’s history. Subsequently, for each potential fault (e.g., Pauli or measurement errors) that might have occured, a (hyper)edge is drawn between the detectors whose signs are flipped by this fault. Each (hyper)edge is assigned a weight based on the probability of the corresponding error occurring [54]. The codes we will be examining are amenable to minimum-weight perfect matching decoding [55], upon decomposing hyperedges into edges. Given a “syndrome” (a set of detectors whose measurements return $-1$ rather than $+1$), the decoder attempts to pair up the triggered detectors to determine a highest-probability correction operation. The decoder succeeds if the error combined with the correction is a trivial logical operator.
Having discussed the general idea of Floquet codes, we now briefly review the commonly studied examples of the code: two variants of the honeycomb code [6, 8], whose performance will also be investigated under biased noise, and the CSS Floquet code [18, 13]. We will later modify the CSS Floquet code to achieve the bias-tailored X3Z3 Floquet code. More details of the honeycomb and CSS Floquet codes are presented in Appendix A.
II.2 Honeycomb codes
We begin by first discussing the honeycomb codes. The first variant is due to Gidney et al. [8], which we call the P6 Floquet code, since its plaquette operators are six-body operators of the form $P^{\otimes 6}$ for $P=X,Y,Z$. We define edge operators of three types: on red edges we define an $XX$ operator, on green edges a $YY$ operator and on blue edges a $ZZ$ operator. We measure edge operators in the periodic sequence $r\rightarrow g\rightarrow b$. When this code is defined with periodic boundary conditions it stores two logical qubits (it is equivalent to the toric code concatenated with a two-qubit repetition code at each time step [6]). The code’s logical operators evolve through the measurement cycle (see Appendix A.1). While the measurement sequence has period 3, the logical operators only return to their initial values (up to signs) with period 6.
We define one stabiliser operator for each plaquette, where red, green, and blue plaquettes host the $X^{\otimes 6}$, $Y^{\otimes 6}$, and $Z^{\otimes 6}$ operators, respectively. These plaquette operators are inferred from edge measurements in two consecutive subrounds. Detectors are formed from consecutive plaquette operator measurements.
The second honeycomb code variant, which we call the XYZ2 honeycomb code, is due to Hastings and Haah [6]. It differs from the P6 code by single-qubit Clifford rotations acting on the qubits. While the XYZ2 code’s edges are still coloured the same as the P6 edges, the XYZ2 code’s edge operators have their Pauli bases defined according to their orientations within each T junction, i.e., horizontal edges are of $Z$ type while the $X$ and $Y$ checks are respectively those edges which are 90 clockwise and counter-clockwise rotated from the horizontal edges (see Appendix A.1). All plaquettes have the same stabiliser operator, i.e., the XYZ2 operator. While the logical operators of the XYZ2 code have the same qubit supports as those of the P6 code, the qubit support bases of the XYZ2 logical operators are not uniform throughout, but involve $X$, $Y$, and $Z$ Paulis [6]. We note that even though these honeycomb code variants have been studied by several works in the literature (e,g., Refs. [6, 8, 7, 9]), there have not been any studies comparing the performance of these two codes under biased noise.
II.3 CSS Floquet code
Having discussed the honeycomb codes, we now give a brief review of another type of Floquet code, the CSS Floquet code, which was first presented in Refs. [18, 13]. We refer the readers to Appendix A.2 for more details on the code. The CSS Floquet code is defined on the same honeycomb lattice. We show the code’s measurement cycle in Fig. 1(a). We measure operators defined on edges using the $r\rightarrow g\rightarrow b$ cycle, but alternate between measuring $XX$ and $ZZ$ operators on these edges. Hence, this code has a period-6 measurement cycle, and its logical operator evolutions also have period-6 (see Appendix A.2 for an illustration). Even though a honeycomb code’s measurement schedule has only a period of 3, for consistency, we will define 1 QEC round to be 6 measurement subrounds for all codes studied in this paper. The check operators, stabilisers and logicals all are either $X$- or $Z$-type and, for this reason, the code is called the CSS Floquet code.
Unlike the honeycomb codes, the CSS Floquet code has no persistent stabiliser operators. Instead, each plaquette stabiliser is repeatedly reinitialised and measured out, with subsequent measurement values compared to form a detector. In contrast to the honeycomb codes where the values of the plaquette stabilisers are inferred from measurements of check operators in two consecutive subrounds, here the plaquette operator measurements take only a single subround [13, 18]. Because of these single-step measurements, at each measurement subround, there is always one type of plaquette operator that anti-commutes with the check measurements, and hence their values are undetermined (see Appendix A.2). These “inactive” plaquettes, however, will be reinitialized at the next measurement subround. As a result, at each time step, the ISG contains $X$-type and $Z$-type stabiliser operators defined on two of the three colours of plaquette and either an $X$-type or a $Z$-type operator defined on the other colour. For instance, after measuring the red-$X$ checks, the ISG contains $X^{\otimes 6}$ and $Z^{\otimes 6}$ blue and green plaquette operators, but only $X^{\otimes 6}$ red plaquette operators, since $Z^{\otimes 6}$ red plaquette operators anti-commute with the red-$X$ checks (see Appendix A.2 for more details).
The CSS Floquet code is naturally suited to a minimum-weight perfect matching (MWPM) decoding, since single-qubit ($X$ or $Z$) Pauli and measurement errors all lead to graph-like syndromes [13]: they each trigger a pair of detectors. There are two decoding graphs formed from the $Z$-type and $X$-type detectors. Only $Y$ errors form hyperedges that need to be decomposed into edges in the two detector graphs.
Figure 2: Persistent symmetry of the X3Z3 Floquet code leads to improved performance under $Z$-biased noise. (a) An example of such a symmetry along a single unshaded strip of the lattice, together with a Pauli-$Z$ error, is shown throughout the measurement cycle. Arrows indicate the type of check ($c\text{A}$ or $c$B for some colour $c$) measured at each subround. A-type plaquette operators whose product forms the symmetry at that subround are highlighted in darker colours. (b) An example of a symmetry after $r$A checks are measured is the product of red and blue A-type plaquettes shown. Since this product is the identity along the vertical unshaded strip and commutes with $Z$ errors on all qubits, the syndromes/anyons [represented by coloured circles in (a)] must appear in pairs along the vertical strips.
Figure 3: The A-type detector graph of the X3Z3 Floquet code under pure dephasing noise. Only a single vertical strip [e.g., that supporting the symmetry in Fig. 2(b)] is shown. The illustrated part of the graph is disconnected from that of any neighbouring strips. Arrows represent only the A-type measurements. The B-type subrounds are not shown, since they do not influence the detectors in this graph. Each node (representing a detector) is placed in the center of a plaquette and at the subround in which the corresponding detector is formed. For example, the bottom layer corresponds to the blue A-check measurements, at which time the red A-type plaquettes are measured.
III X3Z3 Bias-tailored Floquet Code
We now describe the X3Z3 bias-tailored Floquet code, which is shown in Figs. 1(b) and 1(c). This code is related to the CSS Floquet code by Hadamard gates applied to the qubits in alternating strips [i.e., the grey strips in Fig. 1(b)] along vertical non-trivial cycles of the lattice, and is a Floquetified version of the domain wall colour code [45]. As in the CSS Floquet code, the X3Z3 code also has two types of plaquettes and edges: one originates from the Pauli $X$ and the other from the Pauli $Z$ plaquettes and edges in the CSS code before the Hadamard deformation. We refer to these modified operators as A-type and B-type, respectively; these are shown in Fig. 1(b). The measurement sequence is analogous to the CSS Floquet code sequence: $r\text{A}\rightarrow g\text{B}\rightarrow b\text{A}\rightarrow r\text{B}\rightarrow g\text{A}\rightarrow b\text{B}$, where $c\text{A}$ represents the measurement of A-type check operators along $c$-coloured edges, and similarly for $c\text{B}$. Just as with the CSS Floquet code [13], this code can also be defined on a planar lattice with boundary.
The decoding of the X3Z3 Floquet code under $Z$-biased noise is simplified by the presence of a symmetry in the decoding graphs, which we will explain below. This is a space-time analogue of the symmetries present in bias-tailored static codes such as the XZZX code [41] and the domain wall colour code [45]. In such codes, under the infinitely-biased code-capacity noise model (with only single-qubit $Z$ errors, for example), syndromes are forced to come in pairs along one-dimensional (1D) strips of the lattice. This results from strips of stabilisers multiplying together to the identity along one domain and commuting with all $Z$ errors on all qubits: these are 1D symmetries of the stabiliser code under infinitely biased noise. We will refer to (plaquette) stabilisers in the ISG that are flipped by an error as anyons. In bias-tailored static codes, anyons can propagate within strips but cannot move outside the strip without changing their Pauli type. We will show that there also exists a similar symmetry in the bias-tailored Floquet code.
As in the CSS Floquet code, there are two disjoint decoding graphs for the X3Z3 Floquet code: the A-type and B-type graphs, whose edges correspond to single-qubit Pauli errors. In even measurement subrounds, we perform B-type measurements and form detectors for the B-type graph while in odd measurement subrounds we only form detectors for the A-type graph. Note that under a more complicated noise model, such as one that includes measurement or two-qubit errors, edges can exist between the A- and B-type graphs and they are no longer disjoint.
III.1 Persistent symmetry and two-dimensional decoding under biased noise
There is no constant stabiliser group when viewing the X3Z3 Floquet code as a subsystem code (in a subsystem code, the stabiliser group is defined as the center of the gauge group) [13, 18]. Surprisingly, despite this fact, at every time step there do exist operators that form a symmetry under pure dephasing noise (without measurement errors). An example of such a symmetry is shown in Fig. 2 which is a symmetry on the unshaded strips formed by the A-type plaquettes. A similar symmetry can also be found on the shaded strips which are formed by B-type plaquettes. These symmetries, which are formed by plaquette operators of a single type along vertical strips of the lattice at time step $t$, also commute with all operators in $\mathcal{M}_{t+1}$ (the set of check measurements to be performed in the next subround). To see why the symmetries are formed at each measurement subround, consider one particular time step as shown in Fig. 2(b), i.e., after the red A-type measurements. As depicted, the product of red and blue A-type plaquettes is the identity on all qubits in the unshaded vertical strip, and yields Pauli $Z$ operators on some qubits in the neighbouring shaded vertical strips. At this subround, there is a similar symmetry that has the identity on the shaded strips, which is formed by the product of blue and green B-type plaquettes.
As a result of each of these symmetries, Pauli-$Z$ errors create anyons in pairs along vertical strips. An example of such anyons is shown in each measurement subround of Fig. 2(a). Note that the anyons change plaquette locations between subrounds, but always respect the symmetry. The anyons depicted in Fig. 2 represent detectors that would be triggered if the Pauli-$Z$ error shown occurred at that time step. For example, if the $Z$ error occurs immediately after the red A-check measurements, it will trigger a red A detector after subsequent blue A-check measurements and a blue A detector after green A-check measurements.
Unlike the symmetry in static stabiliser codes, this symmetry does not allow for 1D decoding even when measurement errors are absent. While in static codes, only measurement errors produce “time-like” edges, in Floquet codes, even Pauli errors produce time-like edges (between detectors formed at different times). To demonstrate this, we display in Fig. 3 a portion of the A-type detector graph under infinitely-biased single-qubit noise with no measurement errors. As can be seen, there is a disconnected subgraph defined along one strip of the code lattice. Even considering this simple noise model, the decoding graph in the infinite bias regime is two-dimensional (i.e., the graph is planar). This results from the fact that neighbouring plaquettes are measured at different times. We emphasise that although the above discussion is based on $Z$ errors, the same analysis also holds for $X$ errors because $X$ and $Z$ operators are interchangeable for the X3Z3 code. As a result, the performance of the X3Z3 code under $X$-biased noise is expected to be the same as that under the $Z$-biased noise model.
III.2 Bias-preserving parity-check circuits
Figure 4: Depth-2 check measurement circuits using CNOT gates along with resets and single-qubit measurements in the $Z$, $X$ and $Y$ bases for (a) $Z_{1}Z_{2}$, (b) $X_{1}X_{2}$, and (c) $Z_{1}X_{2}$ checks, respectively.
While two-body measurements are native to certain architectures, for example, Majorana [10, 56, 57] and photonic [58, 17] platforms, most architectures require quantum circuits to carry out such measurements. To maintain the high-performance of bias-tailored codes in these hardwares, the parity-check measurement circuits need to be constructed in a bias-preserving manner, such that they do not propagate frequently-occurring errors to rarely-occurring errors. To this end, we design two-qubit parity check circuits that preserve the $Z$ bias on data qubits. That is, the probability of $X$ and $Y$ errors on data qubits after these circuits remains small (proportional to single-qubit $X$ or $Y$ error probabilities). Our bias-preserving parity check measurement circuits areconstructed by using an ancilla circuit connecting two data qubits, which can be realised even in devices with minimal connectivity, such as the heavy-hexagonal layout [53, 52].
For a circuit to be bias preserving, it can be constructed using only gates that do not change the error type under conjugation. For example, one way to construct bias-preserving circuits is to use only CNOT gates in the measurement circuits as they only propagate errors to others with the same Pauli type. For the X3Z3 code, we need to construct bias-preserving circuits for three different kinds of parity checks: $ZZ$, $XX$ and $ZX$. All these three circuits can be constructed using only CNOT gates as shown in Fig. 4. As depicted, these circuits also include resets and measurements of the ancilla qubit in the Pauli bases, i.e., $Z$, $X$ and $Y$ bases for $ZZ$, $XX$ and $ZX$ checks, respectively. Note that the circuits are still bias-preserving even if the resets/measurements in the $X$ and $Y$ bases are compiled in terms of resets/measurements in the $Z$ basis with additional single-qubit Clifford gates. This is because a $Z$ error on any qubit at every time step in the compiled circuits still propagates to data qubits as a $Z$ error, or as a check operator about to be measured.
The parity-check circuits shown in Fig. 4, however, may not preserve the noise bias if errors happen during the CNOT gates. This is because a $Z$ error occurring on the target qubit during the application of a conventional (non-bias preserving) CNOT gate may result in a combination of $Z$ and $X$ errors on the target qubit after the CNOT gate [59]. One way to remedy this problem is to use bias-preserving CNOT gates [59]. These gates, however, are available only in specific platforms such as cat qubits. Here, we propose a more general approach to constructing $Z$-bias preserving parity check circuits without requiring the bias-preserving gates. While our proposed circuits can preserve only a specific type of noise bias, i.e., the $Z$ bias, these circuits are built using conventional (which may not be bias-preserving) two-qubit gates and thus can be implemented in many different architectures.
Our proposed fully bias-preserving parity-check circuits are depicted in Fig. 5. These circuits are constructed using either CZ gates or CNOT gates targeting ancilla qubits, which only propagate Pauli errors on the ancilla qubits as $Z$ errors on the data qubits. While mid-gate $Z$ errors may result in $X$ errors on the ancilla qubits, such errors do not degrade the bias of errors affecting the data qubits. We note that while the $ZZ$ parity check measurements are performed by reading out the ancilla qubits like in standard syndrome extraction circuits, the $XX$ and $ZX$ parity values are obtained by measuring the data qubits after some two-qubit gates are applied on them. These two qubits gates are required for the parity measurements so as not to reveal each individual data qubit’s state. For example, the two mid-circuit $X$-measurements in Fig. 5(b), are equivalent to reading out the $X_{1}X_{a}$ and $X_{2}X_{a}$ parity values, where $X_{a}$ acts on the ancilla qubit which is prepared in the |$\ket{0}$ state. The final $Z_{a}$ measurement disentangles the ancilla qubit, leaving the data qubits in an eigenstate of $X_{1}X_{2}$. The circuit therefore performs the required $XX$ check measurement.
The mixed-type $Z_{1}X_{2}$ parity circuit can be constructed using CZ gates between the data qubits, as shown in Fig. 5(c). If the connections between the data qubits are not available in the device such as in hardwares with a heavy-hexagonal lattice, these CZ gates have to be implemented using ancilla qubits in-between the data qubits. We show such an implementation of next-nearest-neighbour CZ gates in the right-hand side of Fig. 5(c) (this is adapted from Ref. [52]). In the figure, we also present optional measurements, coloured in grey, which provide “flag information” for detecting $Z$ errors on the measurement ancilla that may have propagated to the data qubits.
We emphasise that, even if the $X$ measurements in the parity-check circuits above are implemented using $Z$ measurements sandwiched by Hadamard gates, the circuits are still bias preserving since a $Z$ error occurring between those Hadamard gates is harmless. This is because a $Z$ error happening before the measurement is immediately absorbed by the $Z$ measurement without flipping its outcome, and a $Z$ error after the measurement is a stabiliser of the state (up to a sign), so does nothing. Moreover, the circuit will not change the noise bias significantly even if the Hadamard gates are noisy. The reason is because in many architectures, the single-qubit gate errors are not the predominant error and are usually much smaller than the two-qubit gate errors [60, 61, 62, 63].
Figure 5: Measurement circuits for (a) $ZZ$, (b) $XX$, and (c) $Z_{1}X_{2}$ checks that preserve the $Z$ noise bias on data qubits. Data qubits are labelled by Data1 and Data2, while the middle qubit in each circuit is a measurement ancilla. In (c), the equality can be checked by noting that the depth-3 circuit before and after the mid-circuit measurement performs a next-nearest-neighbour CZ gate, if the measurement ancilla is initially in the $\ket{+}$ state. The grey measurements can be optionally included to provide flag information.
IV Methods
To study the performance of our proposed X3Z3 code, we perform quantum memory experiment simulations, with varying degrees of noise bias $\eta$, of this code along with three other codes: CSS, P6 and XYZ2 Floquet codes. The noise bias $\eta=p_{Z}/(p_{X}+p_{Y})$ is defined as the ratio of the $Z$-error $p_{Z}$ to other errors where the total physical error rate is $p=p_{X}+p_{Y}+p_{Z}$. As the noise asymmetry increases, the noise bias $\eta$ increases from $\eta=0.5$, which corresponds to fully depolarising noise ($p_{X}=p_{Y}=p_{Z}=p/3$), to $\eta=\infty$, corresponding to a pure $Z$-biased noise ($p_{Z}=p$, and $p_{X}=p_{Y}=0$).
We use two different biased noise models which are the generalisations of the standard code-capacity and standard depolarising entangling measurement (SDEM3) [9, 15] noise models to biased noise channels. The code-capacity noise model is one in which only single-qubit errors are applied to all data qubits at every measurement subround. We note that, since for Floquet codes one QEC round consists of several measurement subrounds, the single-qubit noise channel in the code-capacity model is applied several times in one QEC round instead of just once as is the case with static codes. The code-capacity noise model is often used as a preferred initial noise model to study before going to a more involved model as its simplicity often offers insight into understanding the code performance.
On the other hand, the SDEM3 model is a more elaborate error model involving single-qubit noise channels after every single-qubit gate, measurement and reset, a two-qubit noise channel after every multi-Pauli product (MPP) parity measurement gate, and a classical flip after each measurement. As in Ref. [9], we assume that each of these error channels occurs with a total probability $p$. The SDEM3 noise model is therefore close to standard circuit-level noise and would be a more accurate description of a realistic noise channel, particularly in hardwares with native two-qubit measurements, for example, Majorana [56, 10, 57] and photonic [58, 17] architectures. To take into account biased noise for the MPP gates, we here generalise the SDEM3 noise model in Refs. [9, 15]. The details of the biased code-capacity and SDEM3 noise models can be found in Appendix B.
We simulate the circuits and generate the error syndromes using Stim [64] (example Stim circuits and threshold data are provided at Ref. [65]). To decode the error syndromes, we apply a minimum-weight perfect matching (MWPM) decoder, which is implemented using PyMatching [55]. Our simulations are run for codes with effective distances $d_{\mathrm{eff}}\equiv d/2=4,6,8,10,12$, where the effective distance $d_{\mathrm{eff}}$ is defined as the minimum number of faults under SDEM3 depolarising noise that produce a logical error. This is half of the distance $d$ of the code-capacity noise model. The calculated effective distance differs between noise models since in the SDEM3 noise model, two-qubit errors occurring after an MPP parity gate count as a single fault, while under the code-capacity noise model they would count as two faults. We choose lattices with periodic boundary conditions and sizes $L\times 3L/2$. Here, $L=d=2d_{\mathrm{eff}}$ is the lattice length which varies with respect to the code distance. We simulate the memory experiments for $3d_{\mathrm{eff}}$ QEC rounds, with each round containing 6 subrounds of check measurements.
V Results and Discussion
Figure 6: Thresholds ($p_{\mathrm{th}}$) of different Floquet codes as a function of noise bias $\eta$. Codes studied are X3Z3 (blue), CSS (green), P6 (red) and XYZ2 (orange). Results are computed for two different noise models: (a) code-capacity and (b) SDEM3. (a) Inset: Zoom-in threshold plots for the P6 (red) and XYZ2 (orange) codes. For better visualisation, we fit all curves with quadratic splines.
Figure 7: Sub-threshold logical failure probability $p_{L}$ of different Floquet codes as a function of noise bias $\eta$. Codes studied are X3Z3 (blue), CSS (green), P6 (red) and XYZ2 (orange). Results are computed for two different noise models: (a) code-capacity and (b) SDEM3. Inset: Zoom-in subthreshold performance plots of the P6 (red) and XYZ2 (orange) codes. Results are calculated using (a) lattice size $20\times 30$ ($d=20$) for 30 QEC rounds with $p=0.72\%$ and (b) lattice size $24\times 36$ ($d_{\mathrm{eff}}=12$) for 36 QEC rounds with $p=0.5\%$. Each data point is averaged over $10^{6}-10^{8}$ Monte Carlo shots. For better visualisation, we fit all curves with quadratic splines.
Figure 6 shows the thresholds of all codes for various levels of noise bias $\eta$ calculated for (a) code-capacity and (b) SDEM3 noise models. Each threshold is obtained from the intersection of the logical failure probabilities $p_{L}$ vs physical error rate $p$ curves of different code distances $d_{\mathrm{eff}}=d/2=6,8,10$ and 12. The data of $p_{L}$ as a function of $p$ for each code and various noise-bias strengths, used to obtain the thresholds, are presented in Figs. 12-15 of Appendix C.
Since the code-capacity noise model, which considers only single-qubit noise, is a more benign model than the SDEM3 noise model, which also includes two-qubit and measurement errors, the code performance calculated under the code-capacity noise is better than that of the SDEM3 noise. For both noise models, as shown in Figs. 6(a) and 6(b), the performance of the X3Z3 code becomes increasingly better than those of all other tested codes as the noise bias increases. In particular, as the noise changes from fully depolarising to a pure dephasing type, the X3Z3 Floquet code’s threshold increases from $\approx 1.13\%$ to $\approx 3.09\%$ for the code-capacity noise and from $\approx 0.76\%$ to $\approx 1.08\%$ for the SDEM3 noise model. The threshold therefore increases by a factor of 2.7 and 1.4 for the code capacity and SDEM3 noise models, respectively.
It is interesting to note that for the fully depolarising code-capacity noise, all codes have the same threshold. This is because at every measurement subround under this noise model, all the above codes have two types of errors at each fault location (occurring with a total probability $32p/$ 3) that give rise to edges in the detector graph and one type (occurring with probability $p/3$) that produces hyperedges. For the CSS and X3Z3 codes, these hyperedges result from $Y$ Pauli errors. On the other hand, the hyperedge errors in the P6 and XYZ2 codes are those that anti-commute with the check measurements that occur just before and immediately after the errors [12, 22]. Therefore, the Pauli type of the hyperedge errors for the honeycomb codes varies between measurement subrounds. For example, $Z$ errors create hyperedges only when they occur between the $XX$ and $YY$ checks. In summary, all the above codes have the same performance for fully depolarising code-capacity noise because their (weighted) detector hypergraphs are all equivalent.
As shown in Figs. 6(a) and 6(b), while the CSS Floquet code has the same threshold as the X3Z3 code when the noise is fully depolarising, its threshold decreases with increasing noise bias, where its value is $0.752\%$ and $0.668\%$ at infinite bias for the code-capacity and SDEM3 noise, respectively. This decrease is due to the fact that the CSS Floquet code has pure $X$ and pure $Z$ detectors where in the presence of biased noise, half of the detectors, i.e., those with the same type of the dominant error, will become less useful in detecting the biased errors.
Figures 6(a) and 6(b) show that the thresholds for the honeycomb codes have only minor improvements as the noise bias increases, where the thresholds for both honeycomb codes increase only by $\leq 6\times 10^{-4}$. Specifically, the thresholds of the honeycomb codes at $\eta=\infty$ are only about $1.03-1.08$ times larger than their thresholds at $\eta=0.5$ (where their fully-depolarising noise thresholds are $1.13\%$ and $0.585\%$ for the code-capacity and SDEM3 models, respectively). We note that for the SDEM3 noise, the thresholds of both honeycomb codes are lower than those of the X3Z3 and CSS Floquet codes at all noise biases. This is partly explained by noting that the SDEM3 noise model contains measurement errors which give rise to hyperedges in the decoding hypergraphs of the honeycomb codes [12] but only graph-like edges for the CSS and X3Z3 Floquet codes [13]. As we explain below, these hyperedges degrade the MWPM decoder performance.
We find that there are differences, albeit modest ones, between the performance of the XYZ2 and P6 honeycomb codes under biased noise; these differences have not been pointed out before in the literature. While the XYZ2 honeycomb code has a better performance than the P6 honeycomb code for the code-capacity noise model, we find that its performance is inferior to that of the P6 Floquet code for the SDEM3 noise model. This is surprising, considering the similarity of the former to the bias-tailored XYZ2 static code [46]. However, owing to stabilisers being measured in two measurement subrounds, the XYZ2 Floquet code no longer possesses all the symmetries of the static version making the dynamical code not able to inherit all the benefits of its static counterpart.
The performance difference between the XYZ2 and P6 honeycomb codes can be understood from the distribution of hyperedge-like syndromes (namely, those with four triggered detectors) in the detector hypergraphs under the biased SDEM3 noise model. These hyperedges generally degrade the code’s performance when using a matching decoder since they must be decomposed into edges (see Ref. [12] for a detailed description of the decoding in a hyperbolic version of the honeycomb code). An infinitely biased SDEM3 noise model that includes Pauli noise but no measurement errors can already form hyperedges in the honeycomb codes’ detector hypergraphs. In particular, the P6 code can have hyperedges resulting from $Z$ errors occurring only after the red check measurement subrounds, while the hyperedges in XYZ2 code can result from certain $Z$ errors happening in any subround. While there is no difference in the overall number of hyperedges resulting from single-qubit $Z$ errors in the two hypergraphs, the different arrangement of the hyperedges due to the single-qubit errors in these two codes gives rise to different kinds of syndromes for the two-qubit $ZZ$ errors. Specifically, $ZZ$ errors can lead to hyperedge syndromes in the XYZ2 Floquet code but only edge-like syndromes for the P6 code (see Appendix D for details). As a result, in an infinitely biased regime, where Pauli errors after MPP gates are evenly distributed between $Z_{1}$, $Z_{2}$ and $Z_{1}Z_{2}$, there are more hyperedge syndromes triggered in the XYZ2 code than in the P6 code. This is the reason underlying the difference in the thresholds, as shown in Fig. 6(b), of the two honeycomb codes under the SDEM3 noise. Since honeycomb codes are prone to hyperedge errors, better performance of the honeycomb codes might be expected using a correlated MWPM decoder such as the one used in Ref. [8].
As a result of the shift of the thresholds with the noise bias, the subthreshold performance of the codes changes accordingly. As shown in Figs. 7(a) and 7(b), the subthreshold logical failure probability improves significantly for the X3Z3 code, while it becomes only slightly better for the honeycomb codes and deteriorates for the CSS code.
Figure 8: Exponential suppression of the sub-threshold logical failure probabilities $p_{L}$ with respect to $d$ or $d_{\mathrm{eff}}$ for different Floquet codes: CSS (green), XYZ2 (orange), P6 (red), and X3Z3 (blue). Results are calculated for different noise models: (a,b) code-capacity and (c,d) SDEM3 noise models, with a physical error rate $p=0.55\%$. They are also computed using different bias strengths: (a,c) $\eta=1$ and (b,d) $\eta=99$. All curves can be fitted to an exponential decay function $f\propto\mathrm{exp}(-\gamma d)$ or $f\propto\mathrm{exp}(-\gamma d_{\mathrm{eff}})$ where $\gamma$ depends on the bias strength $\eta$ and is an increasing function of $(p_{\mathrm{th}}-p)$. Each data point is averaged over $10^{5}-10^{9}$ shots.
Besides thresholds, another quantity that is of important is the scaling of the code sub-threshold performance with the code distance. To show this, we plot the sub-threshold logical failure probability as a function of $d$ or $d_{\mathrm{eff}}$ for varying levels of noise bias in Fig. 8. As shown, the logical failure probabilities for all codes decrease exponentially with the code distance, i.e., $p_{L}\propto\mathrm{exp}(-\gamma d)$ or $p_{L}\propto\mathrm{exp}(-\gamma d_{\mathrm{eff}})$. Among all codes presented, the X3Z3 code has the largest logical error suppression rate $\gamma$. At higher noise bias, this error suppression rate becomes significantly larger for the X3Z3 code, moderately increases for the two honeycomb codes and decreases for the CSS Floquet code. The reason is that as the noise bias increases, a fixed subthreshold physical error rate moves relatively with respect to the shifting threshold, so that it becomes much further below the threshold for the X3Z3 Floquet code, moves moderately away from the threshold for the two honeycomb codes and becomes closer to the threshold of the CSS code [see Figs. 6(a) and 6(b)].
We also numerically simulate the performance of the X3Z3 Floquet code with a twisted periodic boundary condition. This twisted code has a pure $Z$-type logical operator with a length that scales quadratically with the code distance of the untwisted code. As a result, we find better thresholds and more favorable subthreshold performance for the twisted code over the untwisted one. While these performance benefits hold for the infinite-bias code-capacity noise, they do not apply to other noise biases and models tested (See Appendix E for details).
The error suppression rate $\gamma$ is related to the error suppression factor $\Lambda$ introduced in Ref. [66] which is defined as the reduction factor in the logical failure probability as the code distance increases by 2. Mathematically, it is given by $\Lambda\equiv p_{L}(d)/p_{L}(d+2)=\mathrm{exp}(2\gamma)$ for the code-capacity noise and similarly for the SDEM3 noise model but with $d$ replaced by $d_{\mathrm{eff}}$. In Table 1 of Appendix C.2 we list the values of $\Lambda$ corresponding to the $\gamma$ values shown in Fig. 8.
VI No-go theorems for 1D decoding graphs of dynamical codes
Figure 9: An illustration of (a) Theorem 1 and (b) Theorem 2. We display examples of the $Z$-detector graphs (edges correspond to $Z$ errors and vertices to detectors) that are forbidden by these theorems, for (a) HDCs and (b) GDCs.
[…]
We will start by formalising the above notions via:
Definition 1.
Let $V$ be a set of detectors for a dynamical code. For every qubit and every time step $(q,t)$, i.e., every fault location in the code, define a hyperedge $e_{(q,t)}=(v_{1},v_{2},\ldots)$, where the $v_{i}\in V$ are detectors that return $-1$ outcomes if a single $Z$ error occurs at the fault location $(q,t)$. The $\mathbf{Z}$-detector hypergraph (ZDH) is defined as $G=(V,E)$, where $E=\bigcup_{q,t}e_{(q,t)}$ is the set of all hyperedges. A $Z$-detector graph (ZDG) is a $Z$-detector hypergraph in which all hyperedges are edges (their size is 2).
Definition 2.
A 1D-decodable $Z$-detector graph is one in which all vertices have neighbourhoods of size no greater than 2.
Finally, we define dynamical codes in the following way:
Definition 3.
A honeycomb dynamical code (HDC), with “duration” $T$ and constant “initial(final)-time boundary offsets” $k_{i}$ ($k_{f}$), is a finite-depth measurement circuit acting on qubits of the honeycomb lattice (without boundary). The sets of measurements in the circuit, $\mathcal{M}_{t}$ ($t\in\{1,\ldots,T\}$), are composed of two-body Pauli measurements along coloured edges (either $r$, $g$, or $b$, for each time step) of the lattice, performed sequentially on a state stabilised by some group $\mathcal{S}$. The HDC obeys the following properties:
  1. (a)
    Overlapping measurements in consecutive time steps anti-commute and are supported on different edges,
  2. (b)
    Detectors (for time steps $t>k_{i}$) are associated with plaquettes in the lattice (they have support only on qubits around their associated plaquette),
  3. (c)
    If a single-qubit or two-qubit error occurring in time steps $k_{i}<t<T-k_{f}$ anti-commutes with future measurements, it is detectable, unless the two-qubit error is the same as an edge operator just measured or to be measured in the next time step.
  4. (d)
    All single-qubit errors in time steps $k_{i}<t<T-k_{f}$ have syndromes of weight $>1$.
Let us first comment on our definition of an HDC. The chosen properties are very natural. The anti-commutation of consecutive measurement subrounds ensures “local reversibility” of the code, so as to preserve the quantum information and its locality  [67] (although note that we do not require the code to have logical qubits in our definition). The requirement that consecutive sets of measurements are supported on different edges is not very restrictive: if two (anti-commuting) measurements act on the same edge in consecutive time steps, we can replace the second with a Clifford gate and commute that to the end of the circuit. This merely changes the bases of subsequent measurements without changing their (anti-)commutation. Since the Clifford gates are unimportant, we can ignore them and hence we end up with a shorter duration $T$. We restrict our attention to errors occurring in the temporal “bulk” of the code, $k_{i}<t<T-k_{f}$. This avoids complications due to errors occurring close to the initial and final-time boundaries where detectors here may be formed in different ways than the bulk detectors. Specifically, the detectors associated with errors at the initial and final time boundaries are obtained by comparing the edge or plaquette measurement values with the initial state and final read-out of the physical qubits, respectively. The exclusion of these final time boundary detectors is also required here since we do not include final read-out measurements in our definition of an HDC. In summary, we consider errors occurring in all “detection cells” [13] that are completed before time step $T$ and begin at time step $t\geq 1$, where a detection cell consists of all the spacetime points at which a detector can identify errors.
We will show that an HDC cannot even have a 1D-decodable ZDG, which implies that its decoding graph cannot be equivalent to a set of the much simpler repetition code’s graph. Note that a repetition code graph is not only 1D-decodable but also has a maximum degree of 2, which means there are no double edges between neighbouring vertices. Here, however, we allow for these double edges in our definition of 1D-decodability, and show that this more general property is also impossible. Such double edges naturally arise in Floquet codes, where they correspond to two-qubit undetectable errors (i.e., edge operators just measured or about to be measured).
Using the properties of the HDC as defined above, we now present the first main theorem of our paper (see Appendix F for a proof):
Theorem 1.
A honeycomb dynamical code with duration more than $33k_{f}+k_{i}+$ 3 cannot have a 1D-decodable $Z$-detector graph, where $k_{i}$ and $k_{f}$ are the initial and final time-boundary offsets.
The above theorem implies that no matter how an HDC is bias-tailored, it likely can never achieve as high a threshold as the repetition code or the bias-tailored stabiliser codes. Indeed, for large $T$ ($T\gg k_{i},k_{f}$), there will be at least approximately $6$ times as many edges as vertices in the ZDG, and there is a subgraph with $3/2$ as many edges as vertices without any pair of vertices sharing more than one edge (see Appendix F). This means the graph must have many cycles with length at least three edges.
It is clear that the codes examined in this paper obey the properties of an HDC. In particular, any Floquet codes related to the CSS Floquet code by Hadamard gates applied to a subset of the qubits (as is the case for the X3Z3 Floquet code) do not have 1D-decodable ZDGs. The reason we have focused on dynamical codes that have ZDGs, e.g., Hadamard-deformed CSS codes, instead of codes with hyperedges in their $Z$-detector hypergraphs, is because their decoding graphs are immediate generalisations of that of the CSS code. The fact that 1D-decodable ZDGs are not possible for an HDC suggests that there are not enough independent symmetries to reduce its decoding graph to the extent possible for stabiliser codes. It is interesting to note that even with hyperedges in the $Z$-detector hypergraph, a stabiliser code can still be bias-tailored [38] such that it possesses large enough symmetries to render its thresholds close to the hashing bound [68, 69]. While such high-thresholds are possible for static codes, we do not believe this would be the case for HDCs.
Finally, we consider more general codes defined on an arbitrary graph $G$.
Definition 4.
A general dynamical code (GDC) is a finite-depth measurement circuit acting on qubits associated with vertices of a graph $G$. The sets of measurements $\mathcal{M}_{t}$ are two-body Pauli measurements associated with edges of $G$ (each qubit is in the support of one measurement in each time step), performed sequentially on a state stabilised by some group $\mathcal{S}$. A GDC obeys properties (a) and (d) of an HDC.
Having defined GDCs, we now provide the second main theorem of our paper:
Theorem 2.
A GDC with a duration at least $22k_{f}+k_{i}+$ 2 (i.e., a constant), where $k_{i}$ and $k_{f}$ are the initial and final time-boundary offsets, respectively, cannot have a $Z$-detector graph equivalent to a collection of disjoint repetition codes.
The proof of the theorem is given in Appendix F. Owing to this theorem, we do not expect to see thresholds in dynamical codes as high as 50%. The reason that static codes perform better than dynamical codes under biased noise is that they obey two separate symmetries under an infinitely-biased code-capacity noise model. There is a spatial symmetry that forces syndromes to appear in pairs along 1D strips of the lattice, but there is also a temporal symmetry: syndromes must appear in pairs within a given time slice. These two symmetries simplify the detector graph into a collection of disjoint repetition code decoding graphs, leading to a high performance (up to 50% threshold) of the static codes under biased noise. On the other hand, even for the case where measurement errors are not considered, GDCs necessarily break the temporal symmetry (for HDCs, this is a consequence of Lemma 1 in Appendix F), leading to edges between vertices in different time slices of the detector graph (see Fig. 3). We thus expect that such codes can only possess a single symmetry under an infinitely-biased code-capacity noise model, as the X3Z3 Floquet code does.
VII Conclusions and outlook
In this paper, we introduce the X3Z3 Floquet code, the first bias-tailored dynamical code based on two-qubit parity check measurements. We show that, despite having no constant stabilisers, the code possesses a persistent symmetry under pure dephasing (or pure bit-flip) noise which allows for a simplified decoding. This results in a substantially improved threshold and sub-threshold performance under biased noise, when compared to other Floquet codes. We demonstrate the enhanced performance through our simulation results obtained from using a fast matching decoder and two error models: a simplistic code-capacity noise model and a noise model approaching realistic circuit-level errors. Besides the superior performance of the X3Z3 Floquet code, our results also show that there are differences, albeit modest, in the threshold and subthreshold performance between the XYZ2 and P6 honeycomb codes under biased noise.
To explain why the X3Z3 Floquet code does not reach the same high performance as the bias-tailored static codes, we prove that a dynamical code on the honeycomb lattice (obeying certain assumptions common to standard Floquet codes) cannot have a so-called 1D-decodable $Z$-detector graph, the crucial requirement for the high performance of static codes. Despite this limitation, the bias-tailored X3Z3 Floquet code has the advantage over its static counterpart in that it requires only lower-weight measurements. Specifically for devices without native two-qubit parity measurements, we devise bias-preserving parity check measurement circuits for any qubit architecture, which allow for high-performance implementation of the code. Our work therefore demonstrates that the X3Z3 Floquet code is a leading quantum error correction code especially for devices with limited connectivity such as the hexagonal and heavy-hexagonal architectures.
We now give several directions for future work. While in this paper we use multi-Pauli product gates for the parity check measurements which are native to certain architectures (for example, Majorana [56, 10, 57] and photonic [58, 17] platforms) we leave for future work the study of the code performance using the bias-preserving parity measurement circuits proposed in Sec. III.2. Furthermore, it would also be interesting to investigate the performance of the X3Z3 Floquet code using a better decoder (such as a tensor network decoder) and to consider ways to analytically derive the best achievable thresholds for the X3Z3 Floquet code.
Author Contributions
FS developed the X3Z3 Floquet code and biased noise models. FS also performed the numerical simulations for all Floquet codes investigated. CM characterised the symmetry and decoding features of the codes tested. In addition, CM developed the bias-tailored circuits and the proofs of the no-go theorems. Both authors contributed to the writing of the manuscript.
Acknowledgements.
We would like to thank our colleagues at Riverlane, especially Alexandra Moylett, Ophelia Crawford, and György Gehér for helpful discussions. We are also grateful to Ophelia Crawford and Hari Krovi for their comments on the manuscript. CM would like to thank Benjamin Brown for pointing out a useful circuit for performing next-nearest-neighbour entangling gates.
Appendix A Details of honeycomb and CSS Floquet codes
A.1 Honeycomb code
We here provide details of the two honeycomb code variants discussed in the main text. The codes’ logical operators evolve through their measurement cycles. In Fig. 10(a), we show one set of logical operators evolving through two measurement cycles. While the measurement sequence has a period of 3, the logical operators only return to their initial values (up to signs) with a period of 6.
In Fig. 10(b), we show the plaquette and edge operators for both honeycomb code variants. The P6 code [8] has edge operators with Pauli bases defined by the edge colour, while the XYZ2 code [6] has edge operators with bases defined by the edge’s orientation in the T junction, i.e., horizontal for the $Z$ edges, and rotated 90 clockwise and counter-clockwise from the horizontal edges, for the $X$ and $Y$ edges, respectively [see Fig. 10(b)]. The plaquette operators of the P6 code also differ between plaquettes of different colours, while the plaquette operators of the XYZ2 code are all the same. Each plaquette operator is a product of the edge operators along the plaquette boundary. For example, a red plaquette operator is the product of the green and blue edge operators around its boundary. The plaquette operators are measured in two subrounds. For instance, after measuring the $g$ and $b$ checks sequentially, we can learn the value of each red plaquette by multiplying the measurement values of the green and blue checks around its boundary. Unlike the CSS and X3Z3 Floquet codes, the honeycomb code has constant plaquette stabilisers.
Figure 10: The honeycomb code measurement cycle, logical operators, plaquettes and edges. (a) A distance-4 (under code-capacity noise) P6 code [8] is shown throughout its evolution. The type of check measured at each measurement subround is indicated by the arrows and the highlighted/coloured edges in the lattice. One set of anti-commuting logical operators is shown in yellow and brown-coloured paths. Each logical is a product of Pauli operators on the coloured qubits along its cycle, where red, green and blue qubits support $X$, $Y$, and $Z$ Pauli bases, respectively. The other set of logical operators is offset by three measurement subrounds. (b) Top: Red, green and blue edges/plaquettes of the P6 honeycomb code support $X$, $Y$ and $Z$ operators, respectively. Bottom: All plaquettes of the XYZ2 honeycomb code support the same XYZ2 operator, as shown on the left. Edge operators are defined based on their orientation in the T junction (rather than the edge colour), as shown on the right.
A.2 CSS Floquet code
In Fig. 11, we illustrate the measurement cycle of the CSS Floquet code [18, 13] and the resulting evolution of (one set of) its logical operators. Its logical operators are either of $Z$-type or $X$-type for all time, as shown. We show a $Z$-type logical operator defined on a vertical non-trivial cycle of the lattice. However, there exists another $X$-type logical operator (not shown) defined on that same vertical cycle. One can find this operator (and its anti-commuting partner) by noting the symmetry in the code obtained by proceeding forward by three measurement subrounds ($t\mapsto t+3$) and applying a Hadamard gate to all qubits ($X\leftrightarrow Z$).
The plaquette operator members of the ISG at each step of the measurement cycle are indicated in Fig. 11. The value of each plaquette operator is inferred, in one time step, from the product of the measurement values of three edges surrounding the plaquette. At each measurement subround, only one kind [which is of one type ($X$ or $Z$) and one colour] of plaquette is reinitialised and one other kind is re-measured to form a detector. The reinitialised and re-measured plaquettes have the same ($X$ or $Z$) type as the checks performed at that particular subround, but are of different colours from each other and also from the checks [18, 13]. The colours of the reinitiliased and re-measured plaquettes are respectively the same as the colours of the checks performed just before and immediately after the current subround.
To see how detectors are formed in this code, let us follow a single blue $X$ plaquette through the measurement sequence. After red $X$ checks (top left of Fig. 11), the blue $X$ plaquette operator is measured (we infer its outcome by multiplying together the check measurement outcomes around its boundary). This initialises the plaquette. It is only re-measured after the green $X$ check measurements, four subrounds later, at which point a detector is formed. After this, blue $Z$ checks are measured, which anti-commute with the blue $X$ plaquette operator. Hence, its measurement outcome after this subround is random.
Figure 11: Measurement cycle for the CSS Floquet code. $X$ or $Z$ edge operators are measured alternately, with the colour of edge and Pauli type indicated by the labelled arrows. Plaquettes labelled with $X$ or $Z$ labels host only a stabiliser operator of the corresponding type, while other plaquettes host both $X$- and $Z$-type stabilisers. One set of anti-commuting logical operators is shown by yellow and brown paths. Each logical is a product of only one type of Pauli operators, either $X$ or $Z$, represented by the red or blue coloured qubits.
Appendix B Noise models
Figure 12: Logical failure probability $p_{L}$ of the X3Z3 Floquet code as a function of $x=(p-p_{\mathrm{th}})d^{1/\nu}$ or $x=(p-p_{\mathrm{th}})d_{\mathrm{eff}}^{1/\nu}$ for two different noise models: Code capacity (upper panels) and SDEM3 (lower panels), with various bias strengths: (a,e) $\eta=1$, (b,f) $\eta=9$, (c,g) $\eta=99$, and (d,h) $\eta=999$. We fit the results to the function $A+Bx+Cx^{2}$ (blue line). Insets: Logical failure probability $p_{L}$ vs physical error rate $p$; thresholds $p_{\mathrm{th}}$ are shown by the magenta vertical lines. Results for each distance $d$ and physical error rate $p$ are averaged over $10^{6}-10^{7}$ number of shots.
In this paper, we consider two noise models: code-capacity and entangling measurement (SDEM3) noise models. In the code-capacity noise model, we apply single-qubit Pauli noise on all the data qubits independently at every measurement subround. The single-qubit Pauli noise channel is given by
$\mathcal{E}_{1q}(\rho)=(1-p)\rho+p_{X}\hat{X}\rho\hat{X}+p_{Y}\hat{Y}\rho\hat{Y}+p_{Z}\hat{Z}\rho\hat{Z}.$ (2)
Here, $p\equiv p_{X}+p_{Y}+p_{Z}$ is the total error probability. As in the literature, we define the noise bias as $\eta=p_{Z}/(p_{X}+p_{Y})$ and assume $p_{X}=p_{Y}$. Several values of $\eta$ are worth listing:
  1. 1.
    $\eta=0\rightarrow p_{Z}=0$, and $p_{X}=p_{Y}=p/2$,
  2. 2.
    $\eta=0.5\rightarrow p_{X}=p_{Y}=p_{Z}=p/3$,
  3. 3.
    $\eta=\infty\rightarrow p_{Z}=p$, and $p_{X}=p_{Y}=0$.
Besides the code-capacity noise model, in this paper we also use the SDEM3 noise model which assumes noisy multi-Pauli product (MPP) measurement gates. The SDEM3 noise model applies single-qubit noise after every single-qubit gate, reset and measurement, two-qubit noise after each MPP gate, and a random classical flip after every measurement. We assume each of the error channels occurs with a total probability $p$, as in Ref. [9]. To take into account biased noise, we generalise the SDEM3 model for depolarising noise used in Refs. [9, 15]. Here the single-qubit noise channel has the same form as given in Eq. (2) for the code-capacity noise model.
Figure 13: Logical failure probability $p_{L}$ of the CSS Floquet code as a function of $x=(p-p_{\mathrm{th}})d^{1/\nu}$ or $x=(p-p_{\mathrm{th}})d_{\mathrm{eff}}^{1/\nu}$ for two different noise models: Code capacity (upper panels) and SDEM3 (lower panels), with various bias strengths: (a,e) $\eta=1$, (b,f) $\eta=9$, (c,g) $\eta=99$, and (d,h) $\eta=999$. We fit the results to the function $A+Bx+Cx^{2}$ (blue line). Insets: Logical failure probability $p_{L}$ vs physical error rate $p$; thresholds $p_{\mathrm{th}}$ are shown by the magenta vertical lines. Results for each distance $d$ and physical error rate $p$ are averaged over $10^{6}$ number of shots.
On top of the single-qubit errors, this model also has a two-qubit noise channel applied after each of the MPP gates, which is given by
$\mathcal{E}_{2q}(\rho)=(1-p)\rho+\sum_{O\in\{I,X,Y,Z\}^{\otimes 2}\backslash\{I\otimes I\}}p_{O}\hat{O}\rho\hat{O}.$ (3)
To conform with how the bias is defined for the single-qubit noise channel, we also use $\eta$ to characterize the bias of the two-qubit noise channel. Specifically, we define the bias $\eta$ for the two-qubit noise channel such that
  1. 1.
    $\eta=0\rightarrow p_{ZZ}=p_{IZ}=p_{ZI}=0$ and each of the other probabilities is $p/12$,
  2. 2.
    $\eta=0.5\rightarrow$ each of the Pauli errors occurs with $p/15$, and
  3. 3.
    $\eta=\infty\rightarrow p_{ZZ}=p_{IZ}=p_{ZI}=p/3$, and the other probabilities are 0.
Just as for the single-qubit noise, the above definition of $\eta$ ensures that the two-qubit noise at the three special points, $\eta=0$, $\eta=0.5$ and $\eta=\infty$, are $Z$-error free, depolarising, and pure dephasing, respectively. Given that the two-qubit error probabilities at these $\eta$ values must satisfy the conditions above, we define $\zeta\in[0,1]$ and write the two-qubit Pauli error probabilities as
$p_{ZZ}$ $=p_{ZI}=p_{IZ}=\zeta p/3,$
$p_{O}$ $=(1-\zeta)p/12,$
$\qquad\mathrm{for}\,O\in\{I,X,Y,Z\}^{\otimes 2}\backslash\{II,ZI,IZ,ZZ\},$ (4)
with $\eta$ and $\zeta$ related via:
$\zeta=\frac{3}{5}\left(\frac{\eta}{1+\eta}\right)^{2}+\frac{2}{5}\left(\frac{\eta}{1+\eta}\right).$ (5)
Note that Eq. (B) is defined such that the total probability of all the Pauli errors $\sum_{O\in\{I,X,Y,Z\}^{\otimes 2}\backslash\{I\otimes I\}}p_{O}=p$ for all noise biases $\eta$. Apart from the single- and two-qubit Pauli noise channels, we also apply a classical flip of the measurement results with probability $p$ after each of the single and two-qubit measurements. These measurement flips are uncorrelated with the single and two-qubit noise channels.
Appendix C Details of numerical simulations
Figure 14: Logical failure probability $p_{L}$ of the P6 honeycomb code as a function of $x=(p-p_{\mathrm{th}})d^{1/\nu}$ or $x=(p-p_{\mathrm{th}})d_{\mathrm{eff}}^{1/\nu}$ for two different noise models: Code capacity (upper panels) and SDEM3 (lower panels), with various bias strengths: (a,e) $\eta=1$, (b,f) $\eta=9$, (c,g) $\eta=99$, and (d,h) $\eta=999$. We fit the results to the function $A+Bx+Cx^{2}$ (blue line). Insets: Logical failure probability $p_{L}$ vs physical error rate $p$; thresholds $p_{\mathrm{th}}$ are shown by the magenta vertical lines. Results for each distance $d$ and physical error rate $p$ are averaged over $10^{6}$ number of shots.
For each distance $d_{\mathrm{eff}}=d/2$ (where $d_{\mathrm{eff}}$ and $d$ are the minimum number of fault locations required to produce a logical error for the SDEM3 and code-capacity depolarising noise models, respectively), we construct a $2d_{\mathrm{eff}}\times 3d_{\mathrm{eff}}$ honeycomb lattice with periodic boundary conditions. We run quantum memory experiments for $3d_{\mathrm{eff}}$ QEC rounds or $18d_{\mathrm{eff}}$ measurement subrounds (1 QEC round consists of 6 measurement subrounds). The simulations are performed for four different Floquet codes using different values of physical error rates $p$ and various strengths of noise bias $\eta$. Moreover, we consider two different noise models: code-capacity and SDEM3 noise models as detailed in Appendix B.
We choose a single logical qubit for each Floquet code (they all encode two logical qubits on a lattice with periodic boundary conditions) to test for the logical failure probability. Depending on the codes and parameter regimes, we run different numbers of Monte Carlo shots ranging from $10^{5}-10^{10}$ (larger number of shots for longer code distances and smaller physical error rates) for each of the horizontal and vertical logical observables, which give us the horizontal $p_{H}$ and vertical $p_{V}$ logical failure probabilities. We then report the combined logical error probability:
$p_{L}=1-(1-p_{H})(1-p_{V}),$ (6)
which is an estimate of the probability that either a vertical or a horizontal logical error occurs. Equation (6) assumes that the horizontal and vertical logical errors occur independently. For the infinitely biased code-capacity noise model, one of the logical errors which is of pure $Z$-type or pure $X$-type has zero logical error probability, because the pure biased noise cannot form the other logical operator which is of mixed-type and hence cannot flip the pure-type logical operator. As a result, for the infinitely biased code-capacity noise, the maximum combined logical error probability is $p_{L}=0.5$ [Eq. (6)]. This is in contrast to other noise biases and models where the maximum combined logical error probability is $p_{L}=0.75$ since $p_{V},p_{H}\leq 0.5$.
C.1 Thresholds
Figure 15: Logical failure probability $p_{L}$ of the XYZ2 honeycomb code as a function of $x=(p-p_{\mathrm{th}})d^{1/\nu}$ or $x=(p-p_{\mathrm{th}})d_{\mathrm{eff}}^{1/\nu}$ for two different noise models: Code capacity (upper panels) and SDEM3 (lower panels), with various bias strengths: (a,e) $\eta=1$, (b,f) $\eta=9$, (c,g) $\eta=99$, and (d,h) $\eta=999$. We fit the results to the function $A+Bx+Cx^{2}$ (blue line). Insets: Logical failure probability $p_{L}$ vs physical error rate $p$; thresholds $p_{\mathrm{th}}$ are shown by the magenta vertical lines. Results for each distance $d$ and physical error rate $p$ are averaged over $10^{6}$ number of shots.
To extract thresholds for the codes at various noise-bias strengths, we perform a finite-size data collapse [70, 71] by fitting the data to the curve $A+Bx+Cx^{2}$ where $x=(p-p_{\mathrm{th}})d^{1/\nu}$ or $x=(p-p_{\mathrm{th}})d_{\mathrm{eff}}^{1/\nu}$ for the code-capacity and SDEM3 noise models, respectively. Here, $\nu$ is the critical exponent, $p_{\mathrm{th}}$ is the Pauli threshold and $A,B,C$ are the fit parameters. Figures 12, 13, 14 and 15 show the resulting data collapse and close-to-threshold plots for various bias strengths for the X3Z3, CSS, P6 and XYZ2 Floquet codes, respectively.
C.2 Error suppression factor
Code-capacity SDEM3
$\eta=1$ $\eta=99$ $\eta=1$ $\eta=99$
CSS 1.94 1.39 1.74 1.37
XYZ2 2.27 2.28 1.11 1.15
P6 2.31 2.37 1.12 1.24
X3Z3 2.49 4.26 1.94 3.42
Table 1: Error suppression factors $\Lambda$ for different Floquet codes calculated using two different noise models, code-capacity and SDEM3, with varying levels of noise bias $\eta$. Results are calculated with a physical error rate $p=0.55\%$. The lambda factor is related to the error suppresion rate $\gamma$ shown in Fig. 8
In Table 1, we list the values of the error suppression factor $\Lambda$ which are calculated from the error suppression rate $\gamma$ shown in Fig. 8.
Appendix D Hyperedge errors in honeycomb codes under infinite noise bias
Figure 16: Parts of the detector hypergraphs of (a) P6 and (b) XYZ2 honeycomb codes. Coloured horizontal layers indicate type of measurements performed. Coloured nodes are detectors. Hyperedges and edges between detectors correspond to the syndromes of the single-qubit $Z$ errors shown by the uncoloured circles. Solid (dashed) edges correspond to syndromes due to the $Z$ error on the left (right) qubit. The colour of the (hyper)edge denotes the time step during which the $Z$ error occurs, e.g., a red hyperedge is the syndrome for a $Z$ error occurring immediately after red check measurements. Insets show the syndromes for the two-qubit $ZZ$ error occurring on the two qubits indicated (along a green edge) immediately after green check measurements. This syndrome is edge-like (hyperedge-like) for the P6 (XYZ2) code.
Unlike the CSS and X3Z3 Floquet codes, the single-qubit $X$ and $Z$ Pauli errors in both of the honeycomb code variants can produce hyperedge syndromes [12, 22] if they occur at specific measurement subrounds. These hyperedges degrade the performance of a matching decoder as they must be decomposed into edges, thereby losing some information about the correlations between detectors. We show in Fig. 16 an example of such hyperedges which are formed from single-qubit $Z$ errors. Note that for a single-qubit $Z$ error to produce a hyperedge-like syndrome, it must occur between the measurements of the $XX$ and $YY$ edges of the $Z$-error affected qubit.
First, let us consider the P6 honeycomb code. Since in the P6 code, the edge type is determined by its colour, a single-qubit $Z$ error here generates a hyperedge syndrome (four flipped detectors) if it occurs between red ($XX$) and green ($YY$) check measurements [see the hyperedge highlighted in red in Fig. 16(a)] and an edge-like syndrome (two flipped detectors) if it occur at other measurement subrounds [shown by blue/green edges in Fig. 16(a)]. Notice that the product of the two $Z$ errors occurring simultaneously always results in a graph-like syndrome whenever the error occurs [see inset of Fig. 16(a)]. This is the same for all neighbouring $ZZ$ errors occurring along edges in the lattice.
In the XYZ2 honeycomb code, by contrast, hyperedges can be formed from $Z$ errors occurring at every measurement subround since edges of each colour consist of all check types ($XX$, $YY$ and $ZZ$). The hyperedge-forming $Z$ errors, however, need to occur at specific locations such that they happen between the measurements of the $XX$ and $YY$ edges of the noisy qubits. That is, a $Z$ error with coordinates $(x,y,t)$ could form a hyperedge syndrome for any $t$, given suitably chosen $x$ and $y$. Two such $Z$ errors and their syndromes for different time steps are shown in Fig. 16(b). As can be seen, one $Z$ error (the error on the right) produces a hyperedge if it occurs after red check measurements, while the other (the error on the left) does so only after the green measurement subround. As such, a neighbouring $ZZ$ error occurring at a given time always produce a hyperedge syndrome. For example, if the $ZZ$ error shown (supported by a green edge) in Fig. 16(b) occurs after the green MPP measurements, the resulting syndrome will include four detectors [see inset of Fig. 16(b)].
Appendix E X3Z3 Floquet code with twisted boundary conditions
Figure 17: Evolution of logical operators in the X3Z3 Floquet code with a twisted boundary condition. Number-labelled edges at the top and bottom boundaries indicate the twisted boundary condition. The logical operators are shown in yellow and brown non-trivial cycles.
Figure 18: Thresholds ($p_{\mathrm{th}}$) of the X3Z3 Floquet code with twisted (magenta) and untwisted (blue) periodic boundary conditions as a function of noise bias $\eta$ for two different noise models: (a) code-capacity and (b) SDEM3. We show some of the data points used in obtaining the thresholds in Fig. 19. For better visualisation, we fit all curves with quadratic splines.
Figure 19: Logical failure probability $p_{L}$ of the X3Z3 Floquet code with twisted periodic boundary conditions as a function of $x=(p-p_{\mathrm{th}})L^{1/\nu}$. Data shown are for different lattice lengths $L$ (lattice sizes $L\times 3L/2$). The plots are calculated for two different noise models: code capacity (upper panels) and SDEM3 (lower panels), with various bias strengths: (a,e) $\eta=1$, (b,f) $\eta=9$, (c,g) $\eta=999$, and (d,h) $\eta=\infty$. We fit the results to the function $A+Bx+Cx^{2}$ (blue line). Insets: Logical failure probability $p_{L}$ vs physical error rate $p$; thresholds $p_{\mathrm{th}}$ are shown by magenta vertical lines. Results for each lattice length $L$ and physical error rate $p$ are averaged over $10^{6}-10^{7}$ number of shots.
In this appendix, we present simulation results obtained for the X3Z3 code with a twisted boundary condition. The twisted boundary refers to the case where the qubits at the top row of the honeycomb lattice are connected to qubits at the bottom row, shifted by two columns (see Fig. 17). In the figure, we also show two (of the four) logical operators. The yellow logical operator, which consists of only $Z$ operators, has a length $L^{2}$/2 for a lattice with a size $L\times 3L/2$. This is longer than the length of the untwisted code logical operator, which is only $L$ (or $2L$) [cf. Fig. 1(c)]. As a result, the performance of the twisted code near infinite bias is expected to be better than the untwisted code since it will be harder for the pure dephasing errors to form such a long logical operator.
To support this argument, we present the thresholds of such a code in Fig. 18, where we compare them to the thresholds of the untwisted code. Data used to obtain the thresholds are presented in Fig. 19. We find that there is a stronger finite-size effect for the twisted code compared to the untwisted one as can be seen from the finite-size collapse plots. As shown in Fig. 18, the twisted code has better thresholds than the untwisted code only in the regime close to infinite bias. For the small noise bias regime ($\eta\lesssim 10$), the thresholds of the twisted code are roughly similar to those of the untwisted code. However, in the intermediate bias regime, the thresholds of the twisted code are lower than those of the untwisted code.
Besides the thresholds, we also compute the logical failure probability of the twisted X3Z3 Floquet code as a function of the lattice length $L$. As can be seen from Fig. 20, at infinitely biased code-capacity noise the logical failure probability of the untwisted code decays as $p_{L}\propto\mathrm{exp}(-\gamma_{1}L)$, whereas that of the twisted code decays as $p_{L}\propto\mathrm{exp}(-\gamma_{2}L^{2}-\gamma_{1}L)$. However, when the SDEM3 noise model is considered, both the twisted and untwisted code have only an exponential decay. We note that $\gamma_{1}$ and $\gamma_{2}$ become larger for smaller subthreshold physical error rates.
The super-exponential and exponential decay dependences of the twisted X3Z3 code’s logical failure probability with respect to the lattice length $L$ under the code-capacity and SDEM3 noise models, respectively, are due to the fact that the code distances scale quadratically and linearly with the lattice length under these two different noise models. From the “shortest graph-like error” generated via Stim [64], we find that the measurement and Pauli errors in the SDEM3 model can conspire to produce logical error strings with lengths that scale linearly with the lattice length. Using Stim, we also calculate the size of the shortest graph-like error (i.e., the code distance) of the twisted X3Z3 code under the infinitely $Z$-biased noise and find the following relations:
$d_{\mathrm{twist}}$ $=\frac{d_{\mathrm{untwist}}^{2}}{2}=\frac{L^{2}}{2},\quad$ (7a)
$d_{\mathrm{eff,twist}}$ $=2d_{\mathrm{eff,untwist}}-1=L-1,$ (7b)
for the code-capacity and SDEM3 noise models, respectively. Here $d_{\mathrm{twist}}$ and $d_{\mathrm{eff,twist}}$ are the code distances for the twisted X3Z3 code under the code-capacity and SDEM3 noise, respectively. Similarly, $d_{\mathrm{untwist}}$ and $d_{\mathrm{eff,untwist}}$ are the code distances defined for the untwisted code, while $L$ is the length of the honeycomb lattice with size $L\times 3L/2$. We note that for $\eta<\infty$ of both noise models, the size of the shortest graph-like error calculated from Stim is the same for both the untwisted and twisted codes.
In addition to the subthreshold scaling, as can be seen from Fig. 20, for a fixed physical error rate under the infinite-bias code-capacity noise, the logical failure probabilities of the twisted code are much lower than those of the untwisted case while they are roughly similar for the case of SDEM3 noise. The reason is because the infinite-bias code-capacity threshold for the twisted code is higher than that of the twisted code while they are the same for the SDEM3 case, as shown in Fig. 19.
Figure 20: Sub-threshold logical failure probabilities $p_{L}$ vs lattice length $L$ for the untwisted (blue) and twisted (magenta) X3Z3 Floquet codes. Results are calculated for two different noise models: (a) code-capacity and (b) SDEM3 noise at infinite bias $\eta=\infty$. They are also computed with various physical error rates $p$. All curves can be fitted to decay functions $f\propto\mathrm{exp}(-\gamma_{2}L^{2}-\gamma_{1}L)$ or $f\propto\mathrm{exp}(-\gamma_{1}L)$ where $\gamma_{1,2}$ are increasing functions of $(p_{\mathrm{th}}-p)$. Each data point is averaged over $10^{5}-10^{10}$ number of shots.
Appendix F Details of the no-go theorems for 1D decoding of dynamical codes
In this appendix, we provide the proofs for the two theorems in Sec. VI of the main text. We begin by first proving the following lemma which is going to be used in the proof of Theorem 1.
Lemma 1.
In a honeycomb dynamical code, at each time step $t>1$, no detectors can be formed on neighbouring plaquettes.
Proof.
Without loss of generality, suppose neighbouring plaquettes $F_{1}$ and $F_{2}$ are coloured red and blue, respectively. Suppose, for the sake of contradiction, that they both host detectors at a time step $t>1$. Let $\mathcal{D}_{1},\mathcal{D}_{2}\subset\mathcal{M}_{t}$ be the sets of check measurements from $\mathcal{M}_{t}$ whose products around the plaquettes $F_{1}$ and $F_{2}$, respectively, form the detectors at the corresponding plaquettes at time $t$. That is, the measurements in $\mathcal{D}_{j}$ act only on qubits around $F_{j}$. Since at every time step, only edges of one particular colour are being measured and green edges are shared by blue and red plaquettes, the measurement set $\mathcal{M}_{t}$ at time $t$ can contain only measurements on green edges. Otherwise, either $\mathcal{D}_{1}$ or $\mathcal{D}_{2}$ would have to be empty or they would include measurements on qubits not adjacent to $F_{1}$ or $F_{2}$, respectively, which is not allowed by the definition of $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$.
To form detectors at both plaquettes, we require that, for all $M\in\mathcal{M}_{t-1}$, $[M,\prod_{D\in\mathcal{D}_{j}}D]=0$, for $j=1,2$. Otherwise the plaquette operator outcomes would be indeterministic. However, from property (a) of Definition 3 of an HDC (see Sec. VI of the main text), for each $D\in\mathcal{D}_{1}$, there is a measurement in $\mathcal{M}_{t-1}$ that anti-commutes with it. Therefore, $\prod_{D\in\mathcal{D}_{j}}D$ must have even overlap with all edge measurements in $\mathcal{M}_{t-1}$, for $j=1,2$. This condition can only be satisfied for both $j=1,2$ if all $M\in\mathcal{M}_{t-1}$ are also on green edges (if they are supported on red edges, there will be one measurement with odd overlap with each $D\in\mathcal{D}_{1}$, and similarly for $D\in\mathcal{D}_{2}$ if they are supported on blue edges). But now both $\mathcal{M}_{t-1}$ and $\mathcal{M}_{t}$ are made up of measurements on green edges, contradicting property (a) of Definition 3. ∎
We now prove Theorem 1, which we restate here:
Theorem 1.
A honeycomb dynamical code with duration more than $33k_{f}+k_{i}+$ 3 cannot have a 1D-decodable $Z$-detector graph, where $k_{i}$ and $k_{f}$ are the initial and final time-boundary offsets.
Proof.
We take an error occurring in time step $t$ to mean that it occurs immediately after the measurements in $\mathcal{M}_{t}$. Suppose there are $T$ time steps and $n$ qubits in the HDC. Let us consider part of the ZDG that contains vertices corresponding to all detectors formed in time steps $t>k_{i}$, but only those edges corresponding to detectable errors, i.e., those occurring in time steps $k_{i}<t<T-k_{f}$. We will call this subgraph $\overline{\mathrm{ZDG}}$. The number of vertices in this subgraph, corresponding to the number of detectors formed, is at most $n(T-k_{i})/6$, since at each time step no detectors can be formed on neighbouring plaquettes (see Lemma 1), and there are $n/6$ such non-neighbouring plaquettes, which correspond to all plaquettes of a single colour, in the honeycomb lattice without boundary. Meanwhile, $Z$ errors generate $n$ edges between each time step, resulting in a total of $n(T-1-k_{i}-k_{f})$ edges in the $\overline{\mathrm{ZDG}}$.
Let us now create a bipartition of the qubits into sets $C$ and $D$ such that no qubits in $C$ are adjacent to one another, and similarly for qubits in $D$. This is possible since the honeycomb lattice is bipartite. Let us consider the subgraph of $\overline{\mathrm{ZDG}}$, denoted by $\overline{\mathrm{ZDG}}_{C}$, that contains only edges corresponding to $Z$ errors on qubits in $C$. The subgraph $\overline{\mathrm{ZDG}}_{C}$ therefore contains $n(T-1-k_{i}-k_{f})/2$ edges. Suppose two vertices are connected in this graph by more than one edge. In such a scenario, we can choose any two from the collection of edges between these vertices. These two edges correspond to a weight-2 error, i.e, an error with support only on two qubits, in $C$, which is undetectable. This error is undetectable because the combination of the two errors, where each flips the values of the same two detectors, triggers no detector.
Consider one such undetectable weight-2 error, $Z_{i,t}Z_{j,t^{\prime}}$, where $Z_{i,t}$ denotes an error that acts on qubit $i$ at time step $t$ and similarly for $Z_{j,t^{\prime}}$. We will show that for $Z_{i,t}Z_{j,t^{\prime}}$ not to be detectable, the two errors must act on the same qubit ($i=j$) but at adjacent time steps ($t=t^{\prime}\pm 1$). To this end, we begin by ruling out other possibilities. First, we show that if the qubits $i,j\in C$ are different then the error $Z_{i,t}Z_{j,t^{\prime}}$ will anti-commute with some future measurements. This, along with the fact that $Z_{i,t}Z_{j,t}$ is not an operator just measured or about to be measured (since $i$ and $j$ are not connected by an edge), means that the error is detectable by property (c) of Definition 3 of an HDC (see Sec. VI of the main text). We begin by analysing the case where the two errors occur at the same time step ($t=t^{\prime}$). Suppose that this error commutes with all $M\in\mathcal{M}_{t+1}$. Since every measurement in $\mathcal{M}_{t+1}$ must have support only on one qubit in $C$, then the two check measurements in $\mathcal{M}_{t+1}$, one with support on qubit $i$ and the other on qubit $j$, must have Pauli $Z$ operator supports on those respective qubits. Now consider the measurement $M^{\prime}\in\mathcal{M}_{t+2}$ with support on qubit $i$. By property (a) of Definition 3, one measurement from the set $\mathcal{M}_{t+2}$ that has a support on qubit $i$ cannot act as a Pauli $Z$ operator on that qubit, since it has to anti-commute with the previous edge measurement from $\mathcal{M}_{t+1}$ that also has a $Z$ support on the same qubit. Therefore, $M^{\prime}$ will anti-commute with $Z_{i,t}Z_{j,t}$ which, according to property (c) of Definition 3, means that the error $Z_{i,t}Z_{j,t}$ is detectable. Similarly, if the errors occur on different qubits but at different time steps, they must also be detectable. To show this, we can argue using the reasons mentioned above that $Z_{i,t}$ must anti-commute with a future measurement in set $\mathcal{M}_{t+1}$ or $\mathcal{M}_{t+2}$. Since qubits $i$ and $j$ are not connected by an edge in the honeycomb lattice, the same measurement, however, cannot anti-commute with $Z_{j,t^{\prime}}$.
So the only possibility for the weight-2 error to be undetectable is for the two errors to occur on the same qubit, but at different time steps, i.e., $Z_{i,t}Z_{i,t^{\prime}}$ (note that if the two errors on qubit $i$ occur at the same time step then they cancel each other out, which is equivalent to there being no error at all). Suppose, without loss of generality, that $t<t^{\prime}$. For the weight-2 error to be undetectable, the measurements in between them must commute with $Z_{i}$. Since consecutive measurements with support on the same qubit anti-commute, the two errors acting on qubit $i$ must be separated only by one measurement time step where the in-between measurement commutes with the two errors.
The above analysis implies that vertices in $\overline{\mathrm{ZDG}}_{C}$ can be connected by only at most two edges. All such edges would have to correspond to errors on the same qubit and if there were more than two, we could find a pair of edges that correspond to an undetectable error $Z_{i,t}Z_{i,t^{\prime}}$ with $t^{\prime}>t+1$, which we have shown above to be impossible. Therefore, at most half of the edges in $\overline{\mathrm{ZDG}}_{C}$ can be removed without changing which vertices are connected. This means that there is a spanning subgraph of $\overline{\mathrm{ZDG}}_{C}$ with at least $\lceil n(T-1-k_{i}-k_{f})/4\rceil$ edges with no vertices in the subgraph connected by more than one edge.
All things considered, this implies that the full ZDG has a subgraph with $\lceil n(T-1-k_{i}-k_{f})/4\rceil$ edges and a number of vertices $\leq n(T-k_{i})/6$, such that no two vertices are connected by more than one edge. For $T>3k_{f}+k_{i}+3$, there are more edges than vertices in this subgraph, which means that it cannot have all vertices with neighbourhoods of size $\leq 2$. Therefore, some vertices in the full ZDG also must have neighbourhoods with size greater than 2. ∎
We finally provide the proof for Theorem 2:
Theorem 2.
A GDC with a duration at least $22k_{f}+k_{i}+$ 2, where $k_{i}$ and $k_{f}$ are the initial and final time-boundary offsets, respectively, cannot have a $Z$-detector graph equivalent to a collection of disjoint repetition codes.
Proof.
Since there are $n/2$ measurements in a given time step, there can be at most $n/2$ independent detectors formed from such measurements. However, there will in fact be far fewer for a GDC, because each individual measurement anti-commutes with a measurement from the previous time step resulting in a random value for each edge measurement. Therefore, a detector with a deterministic outcome must be formed from at least a product of two non-deterministic measurements within a time step. Suppose the GDC has a $Z$-detector graph (edges in the ZDG correspond to particular $Z$ errors). Again, we consider the subgraph corresponding to $Z$ errors occurring in time steps $k_{i}<t<T-k_{f}$, where $T$ is the GDC’s duration, and detectors formed in time steps $t>k_{i}$. There are $<n(T-k_{i})/2$ vertices and $n(T-1-k_{i}-k_{f})$ edges in this subgraph. Since the number of vertices of a collection of repetition code decoding graphs cannot be less than the number of edges, this means that the ZDG of a GDC cannot be a collection of repetition code graphs for $T\geq 2k_{f}+k_{i}+2$. ∎